OpenStudy (anonymous):

An angle is its own complement. Find the measure of a supplement to this angle

OpenStudy (anonymous):

...

OpenStudy (anonymous):

..?

OpenStudy (anonymous):

Whats the angle?

OpenStudy (anonymous):

angle is complement which is 90

OpenStudy (anonymous):

a angle thats compliment adds to 90 suppliment adds to 180

OpenStudy (anonymous):

wont it be 90?

OpenStudy (anonymous):

Im still not quite understanding what you are trying to ask...

OpenStudy (anonymous):

If an angle is its own complement, that means it has to be equal to its complement. So an the only angle that's equal to its complement is 45 degrees. Or algebraically, if the angle is x, the complement is 90 - x, correct? So x = 90- x, you get x = 45. Anyways, for this problem then, the supplement of the angle must be an angle that adds up to 180. Since we know it's a 45 degree angle, the other angle is 180 - 45 or 135 degrees.

OpenStudy (anonymous):

so the final answer is?

OpenStudy (ddcamp):

135